five years ago, the director was 3 times older than one of the youngest student, now their combined ages amount to 110 years thus, the present age of the director is?

Answer by mananth(16946)

student's age =x
5 years ago
student's age x-5
director's age =3(x-5)
now student's age =x
director's age = 3(x-5)-5

x+3(x-5)+5=110
x+3x-15+5=110
4x-10=110
4x=120
x=30 student's age now
plug x to get director's age